# Hidden zero modes and topology of multiband non-Hermitian systems

### v1

@@ Abstract
In a finite non-Hermitian system, the number of zero modes does not necessarily reflect the topology of the system. This is known as the breakdown of the bulk-boundary correspondence and has lead to misconceptions about the topological protection of edge modes in such systems. Here we show why this breakdown does occur and that it typically results in hidden zero modes, extremely long-lived zero energy excitations, which are only revealed when considering the singular value instead of the eigenvalue spectrum. We point out, furthermore, that in a finite multiband non-Hermitian system with Hamiltonian $H$, one needs to consider also the reflected Hamiltonian $\tilde{H}$, which is in general distinct from the adjoint $H^{\dagger}$, to properly relate the number of protected zeroes to the winding number of $H$.
@@ I Introduction
In quantum physics, we usually consider observables which are Hermitian operators. These operators have a real eigenspectrum, guaranteeing that expectation values are real and time evolution unitary. However, it has been pointed out about 20 years ago that one can replace the condition of Hermiticity by the less stringent condition of space-time reflection (PT) symmetry and, if this symmetry is unbroken, still obtain a real spectrum [1]. Another motivation to study non-Hermitian Hamiltonians comes from Master equations for open quantum systems [2, 3]. If one ignores quantum jumps, which is a reasonable approximation in certain limits, one can rewrite such a Master equation as a non-Hermitian Hamiltonian [4, 5]. This approach has been used in recent times to analyze and interpret experimental results in optical and magnetic systems [6, 7, 8].
@@
From a theoretical perspective, an important step to understand these phenomena is to classify Gaussian non-Hermitian systems [13, 14]. Here, topology plays a crucial role because it is robust against small perturbations. In one-dimensional Hermitian systems, topological order is only possible if the system possesses additional symmetries. In the case of non-spatial symmetries—such as time reversal, particle-hole, and chiral symmetry—this leads to the tenfold classification of symmetry-protected topological (SPT) order [15]. In contrast, one-dimensional non-Hermitian systems can have a non-trivial topology even without additional symmetries because the eigenspectrum is complex and the determinant of the Bloch Hamiltonian $\det h(k):\,[0,2\pi)\to\mathbb{C}$ as function of momentum $k$ can have a non-zero winding number in the complex plane. It has been shown that if such a winding around a reference point $E$ in the complex plane exists for a single-band model then there will be a skin effect for open boundaries [10]. It has therefore been suggested that the standard bulk topological invariants are not useful to define a bulk-boundary correspondence for non-Hermitian systems and that they have to be replaced by other invariants [16, 17] or by an invariant based on a modified Hermitian Bloch Hamiltonian [14]. However, zero energy edge modes in finite non-Hermitian systems are, in general, fragile to perturbations [9, 18] (see also App. A), implying that they often lack topological protection thus making a bulk-boundary correspondence based on such modes questionable. As an alternative, the singular value spectrum has been put forward [19]. The singular values $s_{i}$ of a system with Hamiltonian $H$ are the square roots of the eigenvalues of $H^{\dagger}H$. For a Hermitian system, we therefore have that $s_{i}=|\lambda_{i}|$ where $|\lambda_{i}|$ are the eigenvalues of $H$. In this case, the standard bulk-boundary correspondence also applies to the singular value spectrum. However, for a non-Hermitian $H$ the eigenvalue and the singular value spectrum are different. While the former is unstable to small perturbations, the latter is stable and important properties can be directly inferred from the topological winding number.
@@ II Topology and winding numbers
$H=\sum_{k}\Psi^{\dagger}_{k}h(k)\Psi_{k}$ (1)
where $\Psi_{k}=(c_{k}^{1},c_{k}^{2},\cdots,c_{k}^{N})$ for a unit cell with $N$ elements, and $h(k)$ is the $N\times N$ Bloch Hamiltonian. If $\det(h(k))$ has no zeroes in $k\in[0,2\pi)$ then we can define a winding number by
@@
In the second line, we have assumed that $h(k)$ is diagonalizable with eigenvalues $\lambda_{j}(k)$, $j=1,\cdots,N$. Note that $\lambda_{j}(0)=\lambda_{j}(2\pi)$ is periodic. This implies, in particular, that if all eigenvalues are real then $\mathcal{I}=0$. A generic one-dimensional Hermitian system is thus always topologically trivial. The only way for a one-dimensional Hermitian system to have non-trivial topology is symmetry-protected topological (SPT) order. These symmetries lead to a block structure of $h(k)$ and topology can be defined based on the properties of these blocks. This leads to the tenfold classification scheme for non-spatial symmetries and to topological crystalline orders for spatial symmetries [20, 21, 22, 23]. In contrast, $\mathcal{I}\neq 0$ is a generic property of non-Hermitian systems which does not require the presence of additional symmetries. The bulk-boundary correspondence for SPT phases of Hermitian systems connects the bulk topological invariant with the number of gapless protected edge modes in a system with boundaries. In a non-Hermitian system, on the other hand, which has a non-zero winding number around a reference energy $E$, we have the skin effect [10]. I.e., a macroscopic number of modes become localized in a system with boundaries but they are fragile with respect to small perturbations.
@@ III Topology in one-dimensional semi-infinite systems
with $\text{ker}(H^{\dagger})=\text{ker}(H^{T})$ and $D$ denotes the dimension. For a Hermitian system, we always have $\text{ind}(H)=0$. The fact that the index can be non-zero tells us that the right and left eigenspectra of systems with semi-infinite boundary conditions are in general not the same. Furthermore, Gohberg’s index theorem directly relates the winding number (2) with the index (4) by [9]
@@ III.1 Example 1
To illustrate these points, we consider an example where $h(k)=1+2\text{e}^{-\text{i}k}$ which means that $\mathcal{I}=-\text{ind}(H)=-1$ and the Fourier coefficients are $h_{0}=1$ and $h_{1}=2$ with all the other ones being equal to zero. Coburn’s lemma then implies that $D(\text{ker}(H))=1$ and $D(\text{ker}(H^{\dagger}))=0$, i.e., $H$ has exactly one right zero mode. In this simple case, we can explicitly construct the right zero mode by considering $Hv=0$ with $v=\begin{pmatrix}v_{1}&v_{2}&\dots\end{pmatrix}^{T}$ and $v_{j}\in\mathbb{C}$. This leads to the simple recurrence relation for the vector coefficients $v_{j}+2v_{j+1}=0$. A normalized solution is given by $v_{j}=(-1)^{j-1}\sqrt{3}/2^{j}$ which means that the zero mode is exponentially localized.
@@ IV Finite systems and hidden zero modes
In a finite Hermitian system with non-trivial SPT order and system size $L$, zero modes can typically be easily identified as eigenvalues which scale as $|\lambda|\sim\exp(-L)$. This, however, is in general not the case in a non-Hermitian system. Here, zero modes can be hidden. This can be understood by considering the example discussed above. If we consider $H_{L}v=0$ for a finite matrix $H_{L}$ and vector $v=\begin{pmatrix}v_{1}&v_{2}&\dots&v_{L}\end{pmatrix}^{T}$ then the recurrence relation above is supplemented by the boundary condition $v_{L}=0$. In this case, the only solution is $v\equiv 0$ which is not a proper solution. The protected zero mode, which does exist in the thermodynamic limit, cannot be found by considering the eigenspectrum of a finite system.
More generally, we call a vector $v$ a hidden zero mode in a finite system with Hamiltonian $H_{L}$ if
@@ IV.1 Example 2
and $h_{j}=0$ otherwise. In this case it is easy to see from Fig. 1 that $D(\text{ker}(H))=0$, $D(\text{ker}(\tilde{H}))=2$, $D(\text{ker}(H^{\dagger}))=1$, and $D(\text{ker}(\tilde{H}^{\dagger}))=1$. For the winding numbers it follows from the index theorem (4,5) that $\mathcal{I}=D(\text{ker}(H^{\dagger}))-D(\text{ker}(H))=1$ and ~$\tilde{\mathcal{I}}=D(\text{ker}(\tilde{H}^{\dagger}))-D(\text{ker}(\tilde{H}))=-1$.
@@
This simple example with unidirectional hopping shows that the zero modes of $H$ and $\tilde{H}$ are very different and that, in the non-scalar case, $\tilde{H}\neq H^{T}$. If we consider this Hamiltonian on a finite line then we find that it does have two zero modes which are fully localized at the right boundary. I.e., the zero modes are not hidden in contrast to our example 1. However, only one is topologically protected while the other one can be removed by a small perturbation.
@@ V Singular values and hidden zero modes
The instability of the eigenspectrum in the non-Hermitian case, and the issues of the two distinct boundaries and the hidden zero modes leads to the question how a bulk-boundary correspondence can be properly formulated. Here, the general theory of truncated Toeplitz operators provides an answer [9]. While relatively little is known about the eigenspectrum of such operators, in particular in the non-scalar case, the $K$-splitting theorem directly relates the index of the operator, and thus also its winding number, to the spectrum of singular values. Assume we have a Hamiltonian $h(k)$ where $\det(h(k))$ has no zeroes in $k\in[0,2\pi)$. Then we define
@@
Next, we want to connect the number of singular values $K$ going to zero with the number of zero modes in the eigenspectrum. Consider the singular value decomposition $H_{L}=USV^{\dagger}$ with $U,V$ unitary and $S$ diagonal and positive. It follows that $H^{\dagger}_{L}H_{L}=VS^{2}V^{\dagger}$. Let us denote the column vectors of $V$ by $v_{n}$, those of $U$ by $u_{n}$ and the singular values on the diagonal of $S$ by $s_{n}$. We then obtain
@@ VII Conclusions
To conclude, we have shown why hidden zero modes appear in finite non-Hermitian systems and how a rigorous bulk-boundary correspondence can be established—based on well-known results for truncated Toeplitz operators—relating the winding number with the singular value spectrum. We also pointed out that zero eigenmodes in finite non-Hermitian systems are often not topologically protected and that there are important differences between the scalar case and multiband models. In experiments, we expect hidden zero modes to be almost indistinguishable from true zero energy eigenmodes because they get mapped by the Hamiltonian exponentially close to zero and thus correspond to extremely long-lived states.
@@ Appendix A Other bulk-boundary correspondences
Here we want to consider other examples which have been studied in the literature and which have lead to some misconceptions about what is and what is not topologically protected in a non-Hermitian system.
@@ A.2 Bloch Hamiltonian for open boundaries
This Hamiltonian is chiral and Hermitian and thus will, according to the standard bulk-boundary correspondence [26], have two protected edge modes if $\mathcal{I}_{1}=-\mathcal{I}_{2}=-1$. This is the case if $|\sqrt{t_{1}^{2}-\left(\frac{\gamma}{2}\right)^{2}}|<1$. Clearly, this is very different from the conditions found for a single or two protected edge modes for the original model (16). More specifically, the model (23) predicts that there are two protected edge modes if either (i) $|\gamma|/2<1$ and $-\sqrt{1+\left(\frac{\gamma}{2}\right)^{2}}<t_{1}<\sqrt{1+\left(\frac{\gamma}{2}\right)^{2}}$ or (ii) $|\gamma|/2>1$ and $\sqrt{-1+\left(\frac{\gamma}{2}\right)^{2}}<t_{1}<\sqrt{1+\left(\frac{\gamma}{2}\right)^{2}}$ or $-\sqrt{1+\left(\frac{\gamma}{2}\right)^{2}}<t_{1}<-\sqrt{-1+\left(\frac{\gamma}{2}\right)^{2}}$. For the example $\gamma=4/3$ considered in Fig. 4, the chiral Hermitian Hamiltonian (23) thus has two protected edge modes if $-\sqrt{\frac{13}{9}}<t_{1}<\sqrt{\frac{13}{9}}$. This is indeed also the regime in which the eigenpectrum of the original model for OBC (17) has two zero eigenmodes. However, as we have aready shown in Fig. 4 these two zero modes are, in general, not topologically protected. It it thus not correct to relate the original non-Hermitian Bloch Hamiltonian to a chiral Hermitian one. This is only allowed if $\mathcal{I}_{1}=-\mathcal{I}_{2}$. Topological protection in regimes where $\mathcal{I}_{1}\neq-\mathcal{I}_{2}$ is fundamentally different and cannot be captured by a Hermitian Bloch Hamiltonian. Instead, the topologically protected modes are hidden and only emerge as exact eigenstates in the thermodynamic limit.

### v2

@@ Abstract
In a finite one-dimensional non-Hermitian system, the number of zero modes does not necessarily reflect the topology of the system. This is known as the breakdown of the bulk-boundary correspondence and has lead to misconceptions about the topological protection of edge modes in such systems. Here we show why this breakdown does occur and that it typically results in hidden zero modes, extremely long-lived zero energy excitations, which are only revealed when considering the singular value instead of the eigenvalue spectrum. We point out, furthermore, that in a finite multiband non-Hermitian system with Hamiltonian $H$, one needs to consider also the reflected Hamiltonian $\tilde{H}$, which is in general distinct from the adjoint $H^{\dagger}$, to properly relate the number of protected zeroes to the winding number of $H$.
@@ I Introduction
In quantum physics, we usually consider observables which are Hermitian operators. These operators have a real eigenspectrum, guaranteeing that expectation values are real and time evolution unitary. However, it has been pointed out about 20 years ago that one can replace the condition of Hermiticity by the less stringent condition of space-time reflection (PT) symmetry and, if this symmetry is unbroken, still obtain a real spectrum [1]. Another motivation to study non-Hermitian Hamiltonians comes from Master equations for open quantum systems [2, 3]. If one ignores quantum jumps, which is a reasonable approximation in certain limits, one can rewrite such a Master equation as a non-Hermitian Hamiltonian [4, 5]. This approach has been used in recent times to analyze and interpret experimental results in optical and magnetic systems [6, 7, 8, 9].
@@
From a theoretical perspective, an important step to understand these phenomena is to classify Gaussian non-Hermitian systems [14, 15]. Here, topology plays a crucial role because it is robust against small perturbations. In one-dimensional Hermitian systems, topological order is only possible if the system possesses additional symmetries. In the case of non-spatial symmetries—such as time reversal, particle-hole, and chiral symmetry—this leads to the tenfold classification of symmetry-protected topological (SPT) order [16]. In contrast, one-dimensional non-Hermitian systems can have a non-trivial topology even without additional symmetries because the eigenspectrum is complex and the determinant of the Bloch Hamiltonian $\det h(k):\,[0,2\pi)\to\mathbb{C}$ as function of momentum $k$ can have a non-zero winding number in the complex plane. It has been shown that if such a winding around a reference point $E$ in the complex plane exists for a single-band model then there will be a skin effect for open boundaries [11]. It has therefore been suggested that the standard bulk topological invariants are not useful to define a bulk-boundary correspondence for non-Hermitian systems and that they have to be replaced by other invariants [17, 18] or by an invariant based on a modified Hermitian Bloch Hamiltonian [15]. However, zero energy edge modes in finite non-Hermitian systems are, in general, fragile to perturbations [10, 19] (see also App. A), implying that they often lack topological protection thus making a bulk-boundary correspondence based on such modes questionable. As an alternative, the singular value spectrum has been put forward [20, 21, 22, 23]. The singular values $s_{i}$ of a system with Hamiltonian $H$ are the square roots of the eigenvalues of $H^{\dagger}H$. For a Hermitian system, we therefore have that $s_{i}=|\lambda_{i}|$ where $|\lambda_{i}|$ are the eigenvalues of $H$. In this case, the standard bulk-boundary correspondence also applies to the singular value spectrum. However, for a non-Hermitian $H$ the eigenvalue and the singular value spectrum are different. While the former is unstable to small perturbations, the latter is stable and important properties can be directly inferred from the topological winding number.
@@ II Topology and winding numbers
$\hat{H}=\sum_{k}\Psi^{\dagger}_{k}h(k)\Psi_{k}$ (1)
where $\Psi_{k}=(c_{k}^{1},c_{k}^{2},\cdots,c_{k}^{N})$ for a unit cell with $N$ elements, and $h(k)$ is the $N\times N$ Bloch Hamiltonian. If $\det(h(k))$ has no zeroes in $k\in[0,2\pi)$—which is also sometimes denoted as $\hat{H}$ having a point gap at $E=0$— then we can define a winding number w.r.t. $E=0$ by
@@
In the second line, we have assumed that $h(k)$ is diagonalizable with eigenvalues $\lambda_{j}(k)$, $j=1,\cdots,N$. Note that $\lambda_{j}(0)=\lambda_{j}(2\pi)$ is periodic. This implies, in particular, that if all eigenvalues are real then $\mathcal{I}=0$. A generic one-dimensional Hermitian system, i.e. a system without additional symmetries, is thus always topologically trivial. The only way for a one-dimensional Hermitian system to have non-trivial topology is symmetry-protected topological (SPT) order. These symmetries lead to a block structure of $h(k)$ and topology can be defined based on the properties of these blocks. This leads to the tenfold classification scheme for non-spatial symmetries and to topological crystalline orders for spatial symmetries [24, 25, 26, 27]. In contrast, $\mathcal{I}\neq 0$ is a generic property of non-Hermitian systems which does not require the presence of additional symmetries. The bulk-boundary correspondence for SPT phases of Hermitian systems connects the bulk topological invariant with the number of gapless protected edge modes in a system with boundaries. In a non-Hermitian system, on the other hand, which has a non-zero winding number around a reference energy $E$, we have the skin effect [11]. I.e., a macroscopic number of modes become localized in a system with boundaries but they are fragile with respect to small perturbations.
@@ III Topology in one-dimensional semi-infinite systems
with $\text{ker}(H^{\dagger})=\text{ker}(H^{T})$ and $D$ denotes the dimension. Here $\text{ker}(H)=\{v|Hv=0\}$ is the kernel of $H$. For a Hermitian system, we always have $\text{ind}(H)=0$. The fact that the index can be non-zero tells us that the right and left eigenspectra of systems with semi-infinite boundary conditions are in general not the same. Furthermore, Gohberg’s index theorem directly relates the winding number (2) with the index (4) by [10]
@@ III.1 Example 1
To illustrate these points, we consider an example where $h(k)=1+2\text{e}^{-\text{i}k}$ which means that $\mathcal{I}=-\text{ind}(H)=-1$ and the Fourier coefficients are $h_{0}=1$ and $h_{1}=2$ with all the other ones being equal to zero. Coburn’s lemma then implies that $D(\text{ker}(H))=1$ and $D(\text{ker}(H^{\dagger}))=0$, i.e., $H$ has exactly one right zero mode. In this simple case, we can explicitly construct the right zero mode by considering $Hv=0$ with $v=\begin{pmatrix}v_{1}&v_{2}&\dots\end{pmatrix}^{T}$ and $v_{j}\in\mathbb{C}$. This leads to the recurrence relation for the vector coefficients $v_{j}+2v_{j+1}=0$. A normalized solution is given by $v_{j}=(-1)^{j-1}\sqrt{3}/2^{j}$ which means that the zero mode is exponentially localized.
@@ IV Finite systems and hidden zero modes
In a finite Hermitian system with non-trivial SPT order and system size $L$, zero modes can typically be easily identified as eigenvalues which scale as $|\lambda|\sim\exp(-L)$. This, however, is in general not the case in a non-Hermitian system. Here, zero modes can be hidden. This can be understood by considering the example discussed above. If we consider $H_{L}v=0$ for a finite matrix $H_{L}$ and vector $v=\begin{pmatrix}v_{1}&v_{2}&\dots&v_{L}\end{pmatrix}^{T}$ then the recurrence relation above is supplemented by the boundary condition $v_{L}=0$. In this case, the only solution is $v\equiv 0$ which is not a proper solution. The protected zero mode, which does exist in the thermodynamic limit, cannot be found by considering the eigenspectrum of a finite system. The only eigenstate for a finite system is $v=\begin{pmatrix}1&0&\dots 0\end{pmatrix}^{T}$ with eigenvalue $E=1$.
More generally, we call a vector $v$ a hidden zero mode in a finite system with Hamiltonian $H_{L}$ if $H_{L}v\neq Ev$ but
@@ IV.1 Example 2
and $h_{j}=0$ otherwise. In this case with unidirectional hopping one obtains directly from $H$ that $D(\text{ker}(H))=0$ and $D(\text{ker}(\tilde{H}))=2$ with the two zero modes in the latter case being $\begin{pmatrix}1&0&\dots&0\end{pmatrix}^{T}$ and $\begin{pmatrix}0&0&1&\dots&0\end{pmatrix}^{T}$. Similarly, $D(\text{ker}(H^{\dagger}))=1$ with zero mode $\begin{pmatrix}1&0&-1&\dots&0\end{pmatrix}^{T}$ and $D(\text{ker}(\tilde{H}^{\dagger}))=1$ with zero mode $\begin{pmatrix}0&1&\dots&0\end{pmatrix}^{T}$. This is illustrated in Fig. 1. This simple example shows that the zero modes of $H$ and $\tilde{H}$ are very different and that, in the non-scalar case, $\tilde{H}\neq H^{T}$. For the winding numbers it follows from the index theorem (4,5) that $\mathcal{I}=D(\text{ker}(H^{\dagger}))-D(\text{ker}(H))=1$ and ~$\tilde{\mathcal{I}}=D(\text{ker}(\tilde{H}^{\dagger}))-D(\text{ker}(\tilde{H}))=-1$.
@@
If we consider this Hamiltonian on a finite line then we find two exact zero modes which are localized at the right boundary. However, adding a small perturbation one zero mode can be removed and the other one becomes a hidden zero mode.
@@ V Singular values and hidden zero modes
The instability of the eigenspectrum in the non-Hermitian case, and the issues of the two distinct boundaries and the hidden zero modes leads to the question how a bulk-boundary correspondence can be properly formulated. Here, the general theory of truncated Toeplitz operators provides an answer [10, 28, 29]. While relatively little is known about the eigenspectrum of such operators, in particular in the non-scalar case, the $K$-splitting theorem [30] directly relates the index of the operator, and thus also its winding number, to the spectrum of singular values. Assume we have a Hamiltonian $h(k)$ where $\det(h(k))$ has no zeroes in $k\in[0,2\pi)$. Then we define
@@
Next, we want to connect the number of singular values $K$ going to zero with the total number of exact and hidden zero modes in the eigenspectrum. Consider the singular value decomposition $H_{L}=USV^{\dagger}$ with $U,V$ unitary and $S$ diagonal and positive. It follows that $H^{\dagger}_{L}H_{L}=VS^{2}V^{\dagger}$. Let us denote the column vectors of $V$ by $v_{n}$, those of $U$ by $u_{n}$ and the singular values on the diagonal of $S$ by $s_{n}$. We then obtain
@@ VII Conclusions
To conclude, we have shown why hidden zero modes appear in finite non-Hermitian systems and how a rigorous bulk-boundary correspondence can be established—based on well-known results for truncated Toeplitz operators—relating the winding number with the singular value spectrum. We also pointed out that zero eigenmodes in finite non-Hermitian systems are often not topologically protected, an issue which we discuss further in App. A, and that there are important differences between the scalar case and multiband models. In experiments, hidden zero modes reveal themselves as extremely long-lived states which can be detected, for example, by preparing the system in this state and let it time evolve [33], see also App. B.
Note added: After conclusion of this work, we became aware of Ref. [34] where the importance of the reflected Hamiltonian $\tilde{H}$ in the non-scalar case was pointed out as well.
@@ Appendix A Other bulk-boundary correspondences
In this Appendix, we want to consider other examples which have been studied in the literature and which have lead to some misconceptions about what is and what is not topologically protected in a non-Hermitian system.
@@ A.2 Bloch Hamiltonian for open boundaries
This Hamiltonian is chiral and Hermitian and thus will, according to the standard bulk-boundary correspondence [35], have two protected edge modes if $\mathcal{I}_{1}=-\mathcal{I}_{2}=-1$. This is the case if $|\sqrt{t_{1}^{2}-\left(\frac{\gamma}{2}\right)^{2}}|<1$. Clearly, this is very different from the conditions found for a single or two protected edge modes for the original model (16). More specifically, the model (23) predicts that there are two protected edge modes if either (i) $|\gamma|/2<1$ and $-\sqrt{1+\left(\frac{\gamma}{2}\right)^{2}}<t_{1}<\sqrt{1+\left(\frac{\gamma}{2}\right)^{2}}$ or (ii) $|\gamma|/2>1$ and $\sqrt{-1+\left(\frac{\gamma}{2}\right)^{2}}<t_{1}<\sqrt{1+\left(\frac{\gamma}{2}\right)^{2}}$ or $-\sqrt{1+\left(\frac{\gamma}{2}\right)^{2}}<t_{1}<-\sqrt{-1+\left(\frac{\gamma}{2}\right)^{2}}$. For the example $\gamma=4/3$ considered in Fig. 4, the chiral Hermitian Hamiltonian (23) thus has two protected edge modes if $-\sqrt{\frac{13}{9}}<t_{1}<\sqrt{\frac{13}{9}}$. This is indeed also the regime in which the eigenpectrum of the original model for OBC (17) has two zero eigenmodes. However, as we have aready shown in Fig. 4 these two zero modes are, in general, not topologically protected. It it thus not correct to relate the original non-Hermitian Bloch Hamiltonian to a chiral Hermitian one. This is only allowed if $\mathcal{I}_{1}=-\mathcal{I}_{2}$ in the original, non-Hermitian model. Topological protection in regimes where $\mathcal{I}_{1}\neq-\mathcal{I}_{2}$ is fundamentally different and cannot be captured by a Hermitian Bloch Hamiltonian. Instead, the topologically protected modes are hidden and only emerge as exact eigenstates in the thermodynamic limit.
Appendix B Experimental detection of hidden zero modes
In this Appendix, we want to describe in more detail how hidden zero modes can be detected experimentally.
For a finite system $H_{L}$, a hidden zero mode $v$ is not an eigenmode, $H_{L}v\neq Ev$, but it does fulfill $||H_{L}v||\to 0$ for $L\to\infty$. Theoretically, we can find these zero modes either by using recurrence relations for the semi-infinite case when $v$ becomes an eigenmode with $Hv=0$ and then truncating the vector $v$ to length $L$ or by using a singular value decomposition. As we have shown in the main text, these two methods are equivalent.
An experimental system will always be finite and a hidden zero mode will be close to but not quite an eigenstate. I.e., a hidden zero mode is a long-lived metastable state. Such states will show up in various response functions—in many-body physics we are used to excitations with a finite lifetime—but the perhaps easiest way to detect them is by a Loschmidt echo type experiment. In the simplest form of such an experiment, the system is prepared in the hidden zero mode $|\Psi\rangle$ and the generalized Loschmidt echo
$\mathcal{L}(t)=\frac{\langle\Psi|\Psi(t)\rangle}{\langle\Psi(t)|\Psi(t)\rangle}$ (24)
is obtained. $\mathcal{L}(t)$ is a measure for how long the time-evolved state $|\Psi(t)\rangle$ remains close to the initial state $|\Psi\rangle$. We call Eq. (24) a generalized Loschmidt echo because the time evolution is not unitary. Note that this implies, in particular that the norm $\langle\Psi(t)|\Psi(t)\rangle$ is not conserved.
Another issue in an experiment is that the non-Hermitian description of an open quantum system is only effective and does ignore quantum jumps. I.e., for this description to work, we also need a postselection where the system is time-evolved many times and only those evolutions are taken into account where the system remains in the considered manifold. We note that exactly this type of experiment has already been performed for a qubit [33]. Here a qubit is prepared in an initial state and the probability monitored that, after postselection, the qubit remains in the initial state under time evolution with an effective non-Hermitian Hamiltonian. Such an experiment, generalized to multiple qubits, would be able to detect hidden zero modes.
To show the expected results, we calculate the Loschmidt echo (24) for example 1 of the main paper as well as for the SSH model studied in the previous section. The results are shown in Fig. 5.
Figure 5: Upper panel: $\mathcal{L}(t)$ for the Hamiltonian in example 1 of the main paper. Lower panel: $\mathcal{L}(t)$ for the SSH chain, (17). In both cases the chain has $L=70$ sites.
For example 1, the Hamiltonian has elements $h_{0}=1$, $h_{1}=2$ and the hidden zero mode is given by $v=\begin{pmatrix}\frac{\sqrt{3}}{2},&-\frac{\sqrt{3}}{4},&\dots,&\frac{\sqrt{3}}{2^{L-1}},&-\frac{\sqrt{3}}{2^{L}}\end{pmatrix}$. If we prepare the system in this state $v$, then $\mbox{Re}[\mathcal{L}(t)]\approx 1$ for times $gt<L$ where $g$ is some effective velocity which depends on the hopping amplitude. In contrast, if we prepare the system in the uniform state $\tilde{v}=\begin{pmatrix}\frac{1}{\sqrt{L}},&\frac{1}{\sqrt{L}},&\dots,&\frac{1}{\sqrt{L}},&\frac{1}{\sqrt{L}}\end{pmatrix}$ then we see a damped oscillation and the Loschmidt echo goes to zero quickly. Also shown in Fig. 5 is the case where the Hamiltonian is detuned by setting $h_{0}=2.2$ but keeping the initial state $v$. Then, there is some immediate decay clearly indicating that $v$ is not a hidden zero mode of the detuned Hamiltonian.
For the SSH model we found that one of the hidden zero modes $v$ in the regime $-1<t_{1}+\frac{\gamma}{2}<1$ is given by Eq. (21). We prepare the system in this state for $t_{1}=-0.2$ and $\gamma=4/3$. Again, the Loschmidt echo shows that this mode is indeed a long-lived metastable state. We also show $\mathcal{L}(t)$ for additional cases. One additional case is for the uniform state $\tilde{v}$ under the same Hamiltonian evolution. The other is for the hidden zero mode $v$ in the presence of a detuned Hamiltonian with the hopping amplitude changed to $t_{1}=-0.1$. In both cases the Loschmidt echo decays quickly.
We conclude that hidden zero modes are important features of non-Hermitian systems and are experimentally observable as long-lived states.